OpenStudy (sloppycanada):

Solve 2x – 5y = 1 by the elimination method. 3x – 4y = -2 x = .66 y = 1

OpenStudy (sepeario):

Do you want me to verify your answers?

OpenStudy (sepeario):

because unfortunate you are wrong.

OpenStudy (sloppycanada):

I was afraid of that...

OpenStudy (sloppycanada):

@Sepeario

OpenStudy (sepeario):

i shall show you by your chosen method of elimination.

OpenStudy (sepeario):

say we have equation 1 (the first) and the second equation equation 2. let's multiply eq 1 by 3 so we have 6x-15y=3 and eq 2 by 2 so we have 6x-8y=-2

OpenStudy (sepeario):

What do you get if you subtract the two newly formed equations?

OpenStudy (sloppycanada):

-23y = 5

OpenStudy (sepeario):

oops the second new equation is 6x-8y=-4

OpenStudy (sepeario):

just check your working, you need to remember that subtracting a negative equals an addition instead.

OpenStudy (sloppycanada):

-7y = 7

OpenStudy (sloppycanada):

y = -1

OpenStudy (sepeario):

good

OpenStudy (sepeario):

now substitute that value to find x.

OpenStudy (sloppycanada):

x = -2

OpenStudy (sepeario):

good job!